# CuriousBot: Interactive Mobile Exploration via Actionable 3D Relational Object Graph

### v1

@@ V CONCLUSION
While our work incorporates diverse manipulation skills for exploration, skill acquisition currently requires experienced roboticists to write and fine-tune the heuristics through trial and error. Developing a scalable skill acquisition process would equip our robot with a broader range of skills and enable more diverse exploration. Additionally, we serialize the object graph for planning, which is adequate in most cases, but it may fall short in scenarios requiring richer semantic details.
Acknowledgement
We would like to thank Jianing Qian, Chao Cao, Yun Chang, Brian Okorn, and Emmanuel Panov for the helpful discussions and feedback on the project. This project is supported in part by the DARPA TIAMAT program (HR0011-24-9-0430) and the Toyota Research Institute (TRI). This article solely reflects the opinions and conclusions of its authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of the sponsors.

### v2

@@ V CONCLUSION
Limitation. While our work incorporates diverse manipulation skills for exploration, skill acquisition currently requires experienced roboticists to write and fine-tune the heuristics through trial and error. These skills might fail if involving heavily occluded scenes. Developing a scalable skill acquisition process would equip our robot with a broader range of skills and enable more diverse exploration. Additionally, since maintaining dynamic scene memory remains an open challenge, we do not update the 3D scene graph after each interaction. In future work, developing a dynamic scene memory that can track objects over time would be highly valuable.
While this work captures various object relations commonly seen in the real world and sufficient for diverse interactive exploration tasks, the real world relations can be even more complex, such as next to. A promising future direction will be capturing object relations automatically using foundation models and develop more expressive representation.